OpenStudy (anonymous):

Please help my bro What is the slope of the line 2x-5y=10 what are the intercepts

hartnn (hartnn):

to find slope,first bring it in the form of y=mx+c try....

OpenStudy (anonymous):

+C ?

hartnn (hartnn):

try to isolate y first...

OpenStudy (anonymous):

If you are given equation directly in the form : \(ax + by = c\) Then slope is given by : \[Slope = -\frac{x \; \; coefficient}{y \; \; coefficient}\]

OpenStudy (anonymous):

Tell me what are x and y coefficients here ??

OpenStudy (anonymous):

2 and 5

OpenStudy (anonymous):

So i divide the two coefficients ?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

what is y coefficient ??

OpenStudy (anonymous):

5 ?

OpenStudy (anonymous):

How ??

OpenStudy (anonymous):

See this like : \[2x + (-5)y = 10\]

OpenStudy (anonymous):

Now tell me what is x and y coefficient here ??

OpenStudy (anonymous):

okay so Y is 2

OpenStudy (anonymous):

Oh My God.. What is the digit attached to x there ??

OpenStudy (anonymous):

\(2x + (-5)y = 10\) Here \(x\) is attached to what ??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

So this is x coefficient and you were right earlier.. And now tell me what is digit attached to y there ??

OpenStudy (anonymous):

can u check me

OpenStudy (anonymous):

Yes show me your work..

OpenStudy (anonymous):

I give you one example: For 5x - 10y = 15 x coefficient is 5 here y coefficient is -10 here and not simply 10.. So now do the same for your question and tell me what are x and y coefficients there ??

OpenStudy (anonymous):